Directions: Solve the following problems stated below. Show your solutions.



a. A -1.49x10^-8 C charge is separated from a -1.15x10^-9 C charge by a distance of 2.01x10^20 m. Determine the repulsive force between the two charges.



b. Two charges (-2.09x10^-18 C and 1.60x10^-19 C) are separated by a distance of 4.01x10^-10 m. Compute the force of attraction between them.

1
Expert's answer
2022-01-24T08:44:40-0500

(a) Given:

"q_1=-1.49*10^{-8}{\rm\: C}, \quad q_2=-1.15*10^{-9}\rm\: C"

"d=2.01*10^{20}\rm\: m"


The Coulomb law says

"F=k\frac{|q_1||q_2|}{d^2}"

"F=9.0*10^9*\frac{1.49*10^{-8}*1.15*10^{-9}}{(2.01*10^{20})^2}"

"F=3.83*10^{-48}\:\rm N"

(b) Given:

"q_1=-2.09*10^{-18}{\rm\: C}, \quad q_2=1.60*10^{-19}\rm\: C"

"d=4.01*10^{-10}\rm\: m"


The Coulomb law says

"F=k\frac{|q_1||q_2|}{d^2}"

"F=9.0*10^9*\frac{2.09*10^{-18}*1.60*10^{-19}}{(4.01*10^{-10})^2}"

"F=1.87*10^{-8}\:\rm N"